OpenStudy (anonymous):

What is the area of a square window that has a perimeter of 20 feet?

jimthompson5910 (jim_thompson5910):

the perimeter is 20 feet so the side length is 20/4 = 5 feet

jimthompson5910 (jim_thompson5910):

the area is therefore 5^2 = 5*5 = 25 square feet

OpenStudy (anonymous):

thank you yur a really big help jim :)

jimthompson5910 (jim_thompson5910):

you're welcome